OpenStudy (antonio_xx2):

I JUST NEED THE ANSWER. What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4?

OpenStudy (anonymous):

Sadly, you can't have just the answer. \[\large\sum_{k=1}^nar^{k-1}=a\frac{1-r^n}{1-r}\] The common ratio is \(r=-4\). The first term occurs when \(k=1\), which would be \[-11=ar^{1-1}~~\iff~~a=-11\] So you have the following series and summ, \[\large\sum_{k=1}^7(-11)(-4)^{k-1}=(-11)\frac{1-(-4)^7}{1-(-4)}\] Well, let me correct myself: you *can* have the answer, just not the simplified version.